# Generative AI Misuse Potential in Cyber Security Education: A Case Study of a UK Degree Program

### v1

@@ IV-C Findings
Three modules were of moderate-to-low exposure, which used lab reports based on practical exercises. Only four assessments in the entire program were deemed to have low exposure ($=$0) of LLM misuse, while six assessments had the highest exposure ($=$5). Concerningly, the modal score was that assessments had very high misuse exposure potential. Some low-exposure assessments were diluted when computing module-level scores. This is because high-exposure assessments sometimes took greater precedence over low-exposure assessments, such as [M10]’s Presentation and [M11]’s Oral Examination, due to the weighting of module assessments. In general, the program contained a disproportionate number of very high-exposure assessments due to its heavy focus on take-home written assessments and project work.
@@ IV-D A Program-level Perspective
The above results focus on building assessment- and module-level exposure scores. We can also consider another metric that evaluates the susceptibility of a program in general to LLM misuse. To this end, we extend the scoring metric in Eq. 1 from [7] to weight modules’ assessment contributions to the program as a whole:
@@
It is important to note that Hickey et al. [7] does not specifically consider other relevant variables in their scoring. For example, the work does not take into account the distribution of cohorts (e.g. home vs. international students), students’ native languages, past experiences of higher education, module delivery modes, and the effects of deadline scheduling. More broadly, they do not consider the availability of academic integrity training, quality of student support services, and access to learning resources.
It is, therefore, reasonable to hypothesise that the program’s actual exposure to LLM misuse may be significantly greater arising from the pressures induced by these additional factors. However, further research is required to understand the precise effects this may have on the tendency for students to (mis)use LLM tools.
@@ V-B Detection Methods
Moreover, an focussing on detecting misuse overzealously may undermine student trust. Students may feel unfairly targeted by automated screening methods, especially in diverse international cohorts where language proficiency varies, and where LLMs may be used legitimately (such as for proofreading). Furthermore, any detection system—however sophisticated—is subject to non-negligible error rates. The resulting interventions may be contentious if students are penalised based on an uncertain evidence base.
@@ V-C Limitations
The approach of weighing programme susceptibility based on credit allocations represents a first attempt at quantifying LLM-exposed programs, but it introduces its own challenges. It assumes that modules’ credit values are equivalent to their contribution to students’ overall learning. Modules with lower credit weights but important practical elements may be disproportionately undervalued, while heavily weighted theoretical modules could skew the overall program exposure metric. The metric also assumes a static model of LLM capabilities and does not account for their evolution. LLMs have already, for example, shown promise in crafting realistic technical presentations and generating real-time verbal responses to questions (e.g. as in a remote oral examination). Ongoing refinement and reflection is thus important in ensuring responsible use of the metrics and results given in this work.

### v2

@@ IV-C Findings
Three modules were of moderate-to-low exposure, which used lab reports based on practical exercises. Only four assessments in the entire program were deemed to have low exposure ($=$0) of LLM misuse, while six assessments had the highest exposure ($=$5). Concerningly, the modal score was that assessments had very high misuse exposure potential. Some low-exposure assessments were diluted when computing module-level scores. This is because high-exposure assessments sometimes took greater precedence over low-exposure assessments, such as [M10]’s Presentation and [M11]’s Oral Examination, due to the weighting of module assessments. In general, the program contained a disproportionate number of highly exposed module due to widespread use of independent report- and project-based assessments.
@@ IV-D A Program-level Perspective
The previous results focussed on assessment- and module-level exposure scores. We may also wish to evaluate the susceptibility of a program at large to LLM misuse. To this end, we extend the scoring metric in Eq. 1 from [7] to weight modules’ assessment contributions to the program as a whole:
@@
It is important to highlight that Hickey et al. [7] do not account for several broader contextual factors in their evaluation of vulnerable assessment types. For instance, their study does not consider cohort composition—such as the proportion of home versus international students—nor does it take into account students’ native languages, prior experiences in higher education, module delivery formats, or the impact of deadline scheduling. Moreover, the analysis overlooks the role of institutional support structures, including academic integrity training, access to learning resources, and student well-being and pastoral services, all of which could influence instances of misuse. Consequently, it is plausible to hypothesise that the program’s actual exposure to LLM misuse may be significantly higher when these additional variables are considered. However, further research is needed to fully understand the extent to which these factors influence students’ propensity to (mis)use LLM tools.
@@ V-B Detection Methods
Moreover, an focussing on detecting misuse overzealously may undermine student trust. Students may feel unfairly targeted by automated screening methods, especially in diverse international cohorts where language proficiency varies, and where LLMs may be used legitimately (such as for proofreading). Furthermore, any detection system—however sophisticated—is subject to non-negligible error rates. The resulting interventions may thus be contentious if students are penalised based on an uncertain evidence base.
@@ V-C Limitations
The approach of weighing programme susceptibility based on credit allocations represents a first attempt at quantifying LLM-exposed programs, but it introduces its own challenges. It assumes that modules’ credit values are equivalent to their contribution to students’ overall learning. Modules with lower credit weights but important practical elements may be disproportionately undervalued, while heavily weighted theoretical modules could skew the overall program exposure metric. The metric also assumes a static model of LLM capabilities and does not account for their evolution. LLMs have already, for example, shown promise in crafting realistic technical presentations [38] and generating real-time verbal responses to questions (e.g. as in a remote oral examination) [39]. Ongoing refinement and reflection is thus important in ensuring responsible use of the metrics and results given in this work.